# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

### v3

@@ 4.1.1 Obstacle avoidance
’Evo-difficult’ Young Spruce forest 2000 42 8/9 0/8
@@ 5.2 Accuracy of forest measurements
The highest estimation accuracies in the literature have been obtained in studies using either considerably heavier and larger ULS systems Hyyppä et al. (2020a, c); Muhojoki et al. (2024) operated in sparse forest environments or in environments with few reference trees Shimabuku, Konoshima, and Ota (2023). The estimation results obtained in this study were close to those obtained in most other studies utilizing camera-based under-canopy drone systems (Kuželka and Surovỳ 2018; He et al. 2025; Krisanski, Taskhiri, and Turner 2020). Furthermore, other studies based on SfM (Shimabuku, Konoshima, and Ota 2023; Kuželka and Surovỳ 2018; He et al. 2025; Krisanski, Taskhiri, and Turner 2020) had more complex flight paths that allowed the full reconstruction most of the tree trunks, whereas in this study, the trunks were only partially reconstructed from image data collected from a single flight line.
Considering the size of the systems, miniaturized lightweight systems were used in the studies by Shimabuku, Konoshima, and Ota (2023), Krisanski, Taskhiri, and Turner (2020), Kuželka and Surovỳ (2018), Prabhu et al. (2024), Cheng et al. (2024), He et al. (2025), and Liang et al. (2024), though the LiDAR-based custom drone systems (Prabhu et al. 2024; Cheng et al. 2024; He et al. 2025; Liang et al. 2024) were still considerably heavier and larger than the camera-based drone systems. From these, LiDAR-based flights in Cheng et al. (2024) and Liang et al. (2024) were the only ones where the systems were reported to fly without commands from the pilot during the flight. Considering the complexity of the environments, the density of the forest in Liang et al. (2024) (1000 trees/ha) was between the forests in this study (650 and 2000 trees/ha). In Cheng et al. (2024), the density of the forest was not reported.
In summary, the approach proposed in this study, camera-based autonomous under-canopy flight and photogrammetric point cloud processing from low-cost onboard stereo camera data, yielded a level of estimation close to the accuracy of most of the studies with SfM drone systems (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; He et al. 2025). In particular, existing studies based on SfM of drones were using commercial drone systems piloted manually, except He et al. (2025) where the camera was added as an additional payload to a custom LiDAR-based drone platform, and the flight mode was not reported. Compared to studies with LiDAR-based miniaturized autonomous drone data collection, the accuracy of the estimation of DBH obtained in this study was higher than in Liang et al. (2024) and close to that in Cheng et al. (2024) even without any GCPs. However, the system utilized in Cheng et al. (2024) has yielded slightly better accuracy with manually piloted flights (Cheng et al. 2024; Prabhu et al. 2024), suggesting that their LiDAR-based system has the potential for higher accuracy of estimation by optimizing autonomous flight trajectories.
@@ 5.3 Potential improvements and future work
The virtual floor and ceiling of Ego-planner-V2 (Zhou et al. 2022) are also defined as static with respect to the takeoff location of the drone, limiting low-altitude flying to relatively flat areas. To capture images at the DBH estimation altitude in uneven forest plots, the flight altitude limits should be modified to be dynamic with respect to the local terrain level.
The drift compensation remains also an important objective for future development. Initial loop closure tests yielded promising results, and future developments could assess the approach using block flights with a lawnmower pattern. In addition to the loop closures, the error in the position estimate could potentially be further reduced by fusing GNSS data to the VIO estimate. Potential methods for GNSS-Fusion include utilizing the global fusion features in VINS-Fusion (Qin et al. 2019a) or replacing VINS-Fusion with another method fusing GNSS-data and VIO, e.g., GVINS (Cao, Lu, and Shen 2022). Although the dense forest canopy blocks the GNSS signals, during long-term operations, the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forest.
@@
In this study, the flight paths were straight flights through the sample plots, and the DBH values were estimated only for the tree trunks recorded during that path. For more comprehensive forest characterization, the trajectory should cover the whole test plot and observe the trees from multiple sides. In the study by He et al. (2025), the DBH estimation error increased significantly when only part of the shape of the tree trunk was observed. This suggests that reconstructing the tree trunk from multiple sides instead of only one side could possibly also reduce the DBH estimate errors obtained with the system utilized in this study. In particular, this is expected to improve the quality of the estimation for trees of large diameter that yielded a higher RMSE in this study than trees with a DBH of less than 30 cm. The global path covering the entire test plot could be implemented with a simple waypoint generation algorithm, such as Liang et al. (2024), or with more advanced exploration methods such as those proposed by Zhou et al. (2021) or Cieslewski, Kaufmann, and Scaramuzza (2017). Another way to achieve area coverage would be utilizing a swarm of drones, as in the original article by Zhou et al. (2022). Furthermore, recent exploration algorithms (Zhou, Xu, and Shen 2023)(Bartolomei, Teixeira, and Chli 2023) are able to utilize drone swarms to further improve and speed up the area coverage. In potential follow-up studies, the DBH estimation potential of the system could also be evaluated in alternative forest environments, for example, in sparser forest environments.
@@
The prototype also opens up possibilities for the study of real-time forest characteristics analysis. In this study, the point clouds were obtained through post-processing. However, Intel RealSense D435 is also capable of generating 3D point clouds in real time. Furthermore, image-based object detection can be performed in real time, for example, with YOLO (Redmon et al. 2016). While only computationally light analysis can be conducted on the onboard computer, real-time analysis with a higher computational burden could be achieved through a remote computer or cloud service via an Internet connection.
Acknowledgements

### v4

@@ 4.1.1 Obstacle avoidance
’Evo-difficult’ Young Spruce forest 2000 42 8/9 0/9
@@ 5.2 Accuracy of forest measurements
The highest estimation accuracies in the literature have been obtained in studies using either considerably heavier and larger ULS systems hyyppa2020under; hyyppa2020comparison; muhojoki24BENCHMARKING operated in sparse forest environments or in environments with few reference trees shimabuku2023diameter. The estimation results obtained in this study were close to those obtained in most other studies utilizing camera-based under-canopy drone systems (kuvzelka2018mapping; he2025estimating360; krisanski2020photodrone). Furthermore, other studies based on SfM (shimabuku2023diameter; kuvzelka2018mapping; he2025estimating360; krisanski2020photodrone) had more complex flight paths that allowed the full reconstruction most of the tree trunks, whereas in this study, the trunks were only partially reconstructed from image data collected from a single flight line.
Considering the size of the systems, miniaturized lightweight systems were used in the studies by shimabuku2023diameter, krisanski2020photodrone, kuvzelka2018mapping, prabhu2024treeFalcon, cheng2024Treescope, he2025estimating360, and liang2024forest, though the LiDAR-based custom drone systems (prabhu2024treeFalcon; cheng2024Treescope; he2025estimating360; liang2024forest) were still considerably heavier and larger than the camera-based drone systems. From these, LiDAR-based flights in cheng2024Treescope and liang2024forest were the only ones where the systems were reported to fly without commands from the pilot during the flight. Considering the complexity of the environments, the density of the forest in liang2024forest (1000 trees/ha) was between the forests in this study (650 and 2000 trees/ha). In cheng2024Treescope, the density of the forest was not reported.
In summary, the approach proposed in this study, camera-based autonomous under-canopy flight and photogrammetric point cloud processing from low-cost onboard stereo camera data, yielded a level of estimation close to the accuracy of most of the studies with SfM drone systems (kuvzelka2018mapping; krisanski2020photodrone; he2025estimating360). In particular, existing studies based on SfM of drones were using commercial drone systems piloted manually, except he2025estimating360 where the camera was added as an additional payload to a custom LiDAR-based drone platform, and the flight mode was not reported. Compared to studies with LiDAR-based miniaturized autonomous drone data collection, the accuracy of the estimation of DBH obtained in this study was higher than in liang2024forest and close to that in cheng2024Treescope even without any GCPs. However, the system utilized in cheng2024Treescope has yielded slightly better accuracy with manually piloted flights (cheng2024Treescope; prabhu2024treeFalcon), suggesting that their LiDAR-based system has the potential for higher accuracy of estimation by optimizing autonomous flight trajectories.
@@ 5.3 Potential improvements and future work
The virtual floor and ceiling of Ego-planner-V2 (zhou2022swarm) are also defined as static with respect to the takeoff location of the drone, limiting low-altitude flying to relatively flat areas. To capture images at the DBH estimation altitude in uneven forest plots, the flight altitude limits should be modified to be dynamic with respect to the local terrain level.
The drift compensation remains also an important objective for future development. Initial loop closure tests yielded promising results, and future developments could assess the approach using block flights with a lawnmower pattern. In addition to the loop closures, the error in the position estimate could potentially be further reduced by fusing GNSS data to the VIO estimate. Potential methods for GNSS-Fusion include utilizing the global fusion features in VINS-Fusion (qin2019global) or replacing VINS-Fusion with another method fusing GNSS-data and VIO, e.g., GVINS (Cao2022GVINS). Although the dense forest canopy blocks the GNSS signals, during long-term operations, the drone could receive corrections to the position estimate in the open areas while relying on the VIO estimate in denser parts of the forest.
@@
In this study, the flight paths were straight flights through the sample plots, and the DBH values were estimated only for the tree trunks recorded during that path. For more comprehensive forest characterization, the trajectory should cover the whole test plot and observe the trees from multiple sides. In the study by he2025estimating360, the DBH estimation error increased significantly when only part of the shape of the tree trunk was observed. This suggests that reconstructing the tree trunk from multiple sides instead of only one side could possibly also reduce the DBH estimate errors obtained with the system utilized in this study. In particular, this is expected to improve the quality of the estimation for trees of large diameter that yielded a higher RMSE in this study than trees with a DBH of less than 30 cm. The global path covering the entire test plot could be implemented with a simple waypoint generation algorithm, such as liang2024forest, or with more advanced exploration methods such as those proposed by Zhou2021FUEL or Cieslewski2017exploration. Another way to achieve area coverage would be utilizing a swarm of drones, as in the original article by zhou2022swarm. Furthermore, recent exploration algorithms (Zhou2023RACER)(Bartolomei2023swarmExploration) are able to utilize drone swarms to further improve and speed up the area coverage. In potential follow-up studies, the DBH estimation potential of the system could also be evaluated in alternative forest environments, for example, in sparser forest environments.
@@
The prototype also opens up possibilities for the study of real-time forest characteristics analysis. In this study, the point clouds were obtained through post-processing. However, Intel RealSense D435 is also capable of generating 3D point clouds in real time. Furthermore, image-based object detection can be performed in real time, for example, with YOLO (Redmon2016yolo). While only computationally light analysis can be conducted on the onboard computer, real-time analysis with a higher computational burden could be achieved through a remote computer or cloud service via an Internet connection.
Funding
@@
Declaration of AI-assisted technologies in the writing process
During the revision of this work, the author(s) used free versions of Grammarly and Writefull (included in the Overleaf LaTeX editor) to improve the writing style in a non-native language manuscript. These tools were only used to detect grammar mistakes and improve the clarity of the text. All suggestions of these writing assistants were manually accepted by the author(s) and no text was autogenerated with generative AI tools. After using these tools, the author(s) reviewed the content and take(s) full responsibility for the content of the publication.